

********************************************************************
UARS WEEKLY STATUS REPORT - 25 FEB 2000
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
SOLSTICE operates normally, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible.
***********
SUSIM
***********
SUSIM continues to gather a near normal set of daily solar scans that
are scheduled for periods of predicted telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM spectrometers continue to be active under PMON control*. AXIS-1
was activated on February 24. At the present time, PEM energy deposition is
not being generated. Recent data configuration changes have affected PEM
software causing multiple failures. These failures are being reviewed by the
PEM team. In order to correct for new telemetry requirements, PEM will most
likely need to update processing code with a patch delivery. At this time,
updated HEPS calibration parameters will be submitted. These will influence
the in-situ energy deposition predictions, so new versions of PEM energy
deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
HALOE continues successful operations with all engineering parameters
within nominal operating range. Except for the first two days after turn on
(February 18, 2000), the tape recorder has captured all the events missed
by the extended TDRS coverage. Efforts by the UARS FOT are greatly
appreciated.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations. The
instrument operations schedule for the 21 March to 25 April
yaw period is being developed.
***********
CDHF
***********
The CDHF continues to notify all PI Software Representatives when Data Gaps
occur in the LEVEL0 data due to the UARS Spacecraft Anomaly.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
35 days WIN (5.11A) PROD (L1-L3) (522R-450C-14F=57L)
4 days SEN (5.0) PROD (1992) (4R-4C=0L)
The CDHF will load December 1999 current day DAAC data in February 2000.
Training for Operations personnel continues on the Rehost systems as
the development progresses.
***********
FOT
***********
REPORTING PERIOD
This report covers 12 February, 2000 (Orbit 46024, GMT Day 043) through 18
February 2000 (Orbit 46128, GMT Day 049).
SPACECRAFT OPERATION
A Reverse-to-Forward Yaw Around maneuver (number 86) was successfully executed
in Orbit 46077 beginning 046/14:25:57 and completing at 046/15:03:07 on 15
February 2000. The SSPP Instruments (ACRIM, SOLSTICE, SUSIM) were enabled as
well as HRDI and PEM in day only operations. Due to the SA position at 269
degrees, the effective night length of the spacecraft after the maneuver was
approximately 48 minutes. Post maneuver analysis of data indicate that during
the Yaw Around, the end-of-night load bus voltage was 24.8 volts and the
average depths of discharge (DODs) on Batteries 2 and 3 were approximately 22.5
and 22.5 percent respectively. The next (forward-to-reverse) YAW is scheduled
for 21 March 2000 (DOY 081).
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
MLS 044/1256 46046 Off
HALOE 045/1200 46061 Off
HALOE 049/1954 46126 On
The beta angle ranged between the angles of 20.6 degrees to 9.4 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
51 degrees on 04 March 1999 (DOY 046).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 11 Febraury 2000 and 18 Febraury 2000 are provided in Appendix D.
The clock error ranged from 8.3 to -2.0 msec during this report period. Clock
rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 01
% ALERTS = 01
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 04
% HALOE = 04
% MLS = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 01
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/021 - 2287.496849
- Transponder B: 00/031 - 2287.498836
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/048 show the observatory data loss to be 565 hours, 38
minutes, 19.273 seconds (an increase of 01 minute 26.880 seconds since the last
report period).
The increase of 1 minutes 26.880 seconds occurred over two (2) days:
- 9.696 seconds on Day 042 due to an IP line hit,
- 01 minute 17.184 seconds on Day 045 due to an IP line hit.
These calculations are under repair. They are being recalculated from the date
where PASS 2 operations were scheduled exclusively.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT finished the algorithms for the Safehold Telemetry Monitors (TMONUs).
Once the uplink tables are completed the FOT will deliver them to the Simulator
support person for implementation and safehold scenario generation.
MEETINGS
None this report period.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for March 21, 2000 (DOY 081). This
maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 18 February 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
46046 044/1256 00-050 MLS Power OFF
46061 045/1200 00-051 HALOE HALOE Stow (Yaw)
46076 046/1301 00-052 OBC YAW AROUND
46115 049/0314 00-053 NBTR NBTR-B MANAGEMENT
46125 049/1843 00-054 NBTR NBTR-B MANAGEMENT
46126 049/1954 00-053 NBTR NBTR-B MANAGEMENT
46126 049/1954 00-055 HALOE Power ON
46127 049/2134 00-056 HALOE HALOE HTR ON
46127 049/2216 00-056 HALOE HALOE HTR OFF
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 18 February 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSPBIMON R-LO PTS Pulse Bus Current Mon 46065
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
MLCLKSTOP MLS Clock Stop Error 46046
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 18 February 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
-None this report period
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 18 February 2000
11 Feb 2000 - GMT Day 042 Beta = 20.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ---------- ----- ------- ------
1 -67.2/-89.6 -3.05/-4.62 +0.0/+0.0 16.0 0.000 0.0
2 +33.6/-22.4 +5.79/+4.83 +31.2/-14.4 25.1 1.041 21.4
3 +89.6/-5.6 +4.51/+3.87 +30.4/-15.2 25.1 1.037 21.3
18 Feb 2000 - GMT Day 049 Beta = 9.4 deg, SA Pos = 269 deg
DIFF V (mV TEMP (C) CURR (am EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- --------- --------- ----------- ------ ------ ------
1 -67.2/-89. -3.68/-4.94 +0.0/+0.0 16.0 0.000 0.0
2 +33.6/-22. +5.15/+3.87 +30.4/-13.6 25.1 1.030 20.2
3 +134.4/11. +3.87/+2.60 +29.6/-15.2 25.0 1.027 19.6
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 18 February 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-004 FOT 46064 045/1623 14170 Generic Late Acq.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiv 512 PB
00-004 FOT see list Feb 00 14170 Generic Late Acq.
SEVNTFW00/07